### v5

# Completions of Kleene’s second model

Abstract.
We investigate completions of partial combinatory algebras (pcas), in particular of Kleene’s second model $\mathcal{K}_{2}$ and generalizations thereof. We consider weak and strong notions of embeddability and completion that have been studied before in the literature. It is known that every countable pca can be weakly embedded into $\mathcal{K}_{2}$, and we generalize this to arbitrary cardinalities by considering generalizations of $\mathcal{K}_{2}$ for larger cardinals. This emphasizes the central role of $\mathcal{K}_{2}$ in the study of pcas. We also show that $\mathcal{K}_{2}$ and its generalizations have strong completions.
Key words and phrases:
partial combinatory algebra, embeddings, completions
2010 Mathematics Subject Classification:
03B40, 03D45, 03D80.
1. Introduction
Combinatory algebra supplies us with a large variety of abstract models of computation. Kleene’s second model $\mathcal{K}_{2}$, first defined in Kleene and Vesley [12], is a partial combinatory algebra defined by an application operator on reals. (This in contrast with Kleene’s first model $\mathcal{K}_{1}$, defined in terms of application on the natural numbers, which is the setting of classical computability theory.) Let $\omega^{\omega}$ denote Baire space, and let $\alpha,\beta\in\omega^{\omega}$. Then the application operator in $\mathcal{K}_{2}$ can be described by
$\alpha\cdot\beta=\Phi_{\alpha(0)}^{\alpha\oplus\beta}.$ (1)
Here $\Phi_{e}$ denotes the $e$-th Turing functional, and $\alpha\oplus\beta$ the join of $\alpha$ and $\beta$. The application is understood to be defined if the right hand side is total. This definition of application in $\mathcal{K}_{2}$ is the one from Shafer and Terwijn [22], which is not the same as the original coding from Kleene but essentially equivalent to it, and more user friendly. The sense in which the two codings are equivalent is explained in Golov and Terwijn [10]. In Section 4, we will also use the original coding of $\mathcal{K}_{2}$ (see Definition 4.1) when we are discussing larger cardinals, for which we do not have a machine model available.
We will also make use of several variants of $\mathcal{K}_{2}$. $\mathcal{K}_{2}$ has an effective version $\mathcal{K}_{2}^{\mathrm{eff}}$, which is obtained by simply restricting $\mathcal{K}_{2}$ to computable elements of $\omega^{\omega}$. The van Oosten model $\mathcal{B}$, first defined in [18], is a variant of $\mathcal{K}_{2}$ obtained by extending the definition of $\mathcal{K}_{2}$ to partial functions. This also has an effective version, denoted by $\mathcal{B}^{\mathrm{eff}}$.
A completion of a pca is an embedding into a (total) combinatory algebra. Here one can study weak and strong kinds of completion, see Definition 2.1. In the strong version, the combinators $k$ and $s$ that every pca possesses are considered as part of the signature, and in the weak version they are not. Both kinds of embedding have been studied in the literature. The question (posed by Barendregt, Mitschke, and Scott) whether every pca has a strong completion was answered in the negative by Klop [13], see also Bethke, Klop, and de Vrijer [5]. On the other hand, $\mathcal{K}_{2}$ does have strong completions, see Theorem 3.3 below. Weak completions and embeddings were studied for example in Bethke [3], Asperti and Ciabattoni [1], Shafer and Terwijn [22], Zoethout [25], Golov and Terwijn [10], and more recently in Fokina and Terwijn [9]. It follows from the main result in Engeler [6] that every pca has a weak completion. Hence fixing the combinators $k$ and $s$ or not in embedding results makes a big difference. There is an even weaker notion of embeddability of pcas introduced by Longley, see [15] and [16], that is useful in the study of realizability. We will not consider this notion here, but simply note that the results about weak completions obtained here are the strongest possible. Extensions of Kleene’s second model containing specific functions were studied in van Oosten and Voorneveld [20].
It is known that every countable pca is weakly embeddable into $\mathcal{K}_{2}$. This is optimal, since this does not hold for strong embeddings. We generalize this to larger cardinalities by considering a generalization $\mathcal{K}_{2}^{\kappa}$ for larger cardinals $\kappa$ (Section 4). We show that $\mathcal{K}_{2}$ and $\mathcal{K}_{2}^{\kappa}$ have strong completions (Theorems 3.3 and 4.4), and prove an embedding result for $\mathcal{K}_{2}^{\kappa}$ (Theorem 5.1).
[…]
2. Preliminaries
A partial combinatory algebra (pca) is a set $\mathcal{A}$ together with a partial map $\cdot$ from $\mathcal{A}\times\mathcal{A}$ to $\mathcal{A}$, that has the property that it is combinatorially complete (see below). We usually write $ab$ instead of $a\cdot b$, and if this is defined we denote this by $ab\!\downarrow$. By convention, application associates to the left, so $abc$ denotes $(ab)c$. We always assume that our pcas are nontrivial, i.e. have more than one element. Combinatorial completeness is equivalent to the existence of elements $k$ and $s$ with the following properties for all $a,b,c\in\mathcal{A}$:
[…]
We have the following notions of embedding for pcas. To distinguish applications in different pcas, we also write $\mathcal{A}\models a\cdot b\!\downarrow$ if this application is defined in $\mathcal{A}$.
Definition 2.1.
For given pcas $\mathcal{A}$ and $\mathcal{B}$, an injection $f:\mathcal{A}\to\mathcal{B}$ is a weak embedding if for all $a,b\in\mathcal{A}$,
$\mathcal{A}\models ab\!\downarrow\;\Longrightarrow\;\mathcal{B}\models f(a)f(b)\!\downarrow\,=f(ab).$ (4)
If $\mathcal{A}$ embeds into $\mathcal{B}$ in this way we write $\mathcal{A}\hookrightarrow\mathcal{B}$. If in addition to (4), for a specific choice of combinators $k$ and $s$ of $\mathcal{A}$, $f(k)$ and $f(s)$ serve as combinators for $\mathcal{B}$, we call $f$ a strong embedding.
A (total) combinatory algebra $\mathcal{B}$ is called a weak completion of $\mathcal{A}$ if there exists an embedding $\mathcal{A}\hookrightarrow\mathcal{B}$. If the embedding is strong, we call $\mathcal{B}$ a strong completion.
There exists pcas that do not have a strong completion, see Klop [13] and Bethke, Klop, and de Vrijer [5]. On the other hand, it follows from Engeler [6] that every pca has a weak completion.111 The main result in [6] is stated only for structures with one binary operator, but on p390 it is mentioned that the representation theorem can be generalized to multiple and also to partial operations, so that in particular it also holds for pcas. This claim is further substantiated in the supplementary note [7]. So in particular we see that weak and strong embeddings and completions are different.
We can see that every countable pca has a weak completion as follows. In Golov and Terwijn [10, Corollary 6.2] it was shown that every countable pca is weakly embeddable into $\mathcal{K}_{2}$. Now $\mathcal{K}_{2}$ is weakly completable, as it weakly embeds, for example, into Scott’s graph model $\mathcal{G}$ (see Scott [21] for the definition, and [10] for the embedding), which is a total combinatory algebra related to enumeration reducibility. Hence every countable pcas has a weak completion. Note that this includes the counterexample for strong completions from Bethke et al. [5], which is countable.
In the next section we show that $\mathcal{K}_{2}$ and $\mathcal{K}_{2}^{\mathrm{eff}}$ have strong completions. Note that none of these completions is computable, as (total) combinatory algebras are never computable. The complexity of completions was studied in [24].
3. Unique head normal forms
In this section we give two different arguments showing that $\mathcal{K}_{2}$ is strongly completable. We start by verifying that $\mathcal{K}_{2}$ satisfies the criterion from Bethke, Klop, and de Vrijer [4] about unique head normal forms (hnfs). This argument that $\mathcal{K}_{2}$ is completable is longer than the short argument at the end of the section, but it is still interesting because it yields as a completion a certain term model that is potentially different, and it also serves as a template for the later generalization to larger cardinalities (Theorem 4.4).
Given a pca $\mathcal{A}$ and a choice of combinators $s$ and $k$, the head normal forms (hnfs) of $\mathcal{A}$ are all elements of the form $s$, $k$, $ka$, $sa$, and $sab$, with $a,b\in\mathcal{A}$. Each of these five types of hnf is called dissimilar from the other four.
Definition 3.1 (Bethke et al. [4]).
$\mathcal{A}$ has unique hnfs if
  • no two dissimilar hnfs can be equal in $\mathcal{A}$,
  • Barendregt’s axiom holds in $\mathcal{A}$: for all $a,a^{\prime},b,b^{\prime}\in\mathcal{A}$,
    $sab=sa^{\prime}b^{\prime}\Longrightarrow a=a^{\prime}\wedge b=b^{\prime}.$ (5)
Bethke et al. [4] confirmed a conjecture from Klop [13], namely that every pca with unique hnfs is strongly completable.
Lemma 3.2.
$\mathcal{K}_{2}$ and $\mathcal{K}_{2}^{\mathrm{eff}}$ have unique hnfs.
Proof.
As mentioned in the introduction, we will use the coding of $\mathcal{K}_{2}$ with application defined by
$\alpha\cdot\beta=\Phi_{\alpha(0)}^{\alpha\oplus\beta}$
for all $\alpha,\beta\in\omega^{\omega}$.
We verify that we can choose $k$ and $s$ in $\mathcal{K}_{2}$ with their defining properties (2) and (3), and such that all hnfs are different. First we note that since $\mathcal{K}_{2}$ is nontrivial (i.e. has more than one element) we will automatically have $s\neq k$ (cf. Barendregt [2, Remark 5.1.11]).
Somewhat abusing terminology (common in set theory), we will refer to the elements of $\omega^{\omega}$ as reals. Without spelling out all syntactic details, we simply note that we can define computable reals $k$ and $s$ such that the following hold:
  • $ka=\Phi_{k(0)}^{k\oplus a}$ is a real that stores $a$, prepended by a code $ka(0)$ that does not need $a$, so $ka(0)=m$ for all $a$.
  • $kab=\Phi_{ka(0)}^{ka\oplus b}=\Phi_{m}^{ka\oplus b}$. Here $m$ is a code that unpacks $a$ from $ka\oplus b$.
  • $sa=\Phi_{s(0)}^{s\oplus a}$ is a real that stores $a$, prepended by a code $sa(0)$ not depending on $a$, so $sa(0)=n$ for all $a$.
  • $sab=\Phi_{sa(0)}^{sa\oplus b}=\Phi_{n}^{sa\oplus b}$ is a real that stores $a$ and $b$, prepended by a code $sab(0)$ that does not need $a$ and $b$, so $sab(0)=p$ for all $a$ and $b$.
  • $sabc=\Phi_{sab(0)}^{sab\oplus c}=\Phi_{p}^{sab\oplus c}$. Here $p$ is a code that unpacks $a$ and $b$ from $sab\oplus c$, and then computes $ac(bc)$.
To turn these informal definitions into formal ones, one would first give more explicit codings of $ka$, $sa$, and $sab$, and then formally define $m,n,p$. However, the details of how one would store these reals are not essential, and the above definitions should suffice to convince the reader of the existence of $k$ and $s$. Note that $k$ and $s$, although elements of $\omega^{\omega}$, are essentially finite objects coding finite programs, hence computable. Also, all hnfs defined above are different. Note that the definition of $sab$ guarantees Barendregt’s axiom (5), since $a$ and $b$ can be extracted from $sab$. This shows that we can choose $k$ and $s$ so that $\mathcal{K}_{2}$ has unique hnfs. The argument for $\mathcal{K}_{2}^{\mathrm{eff}}$ is the same, since we may use the exact same computable reals $k$ and $s$ for $\mathcal{K}_{2}^{\mathrm{eff}}$. So also $\mathcal{K}_{2}^{\mathrm{eff}}$ has unique hnfs. ∎
Theorem 3.3.
$\mathcal{K}_{2}$ and $\mathcal{K}_{2}^{\mathrm{eff}}$ are strongly completable.
Proof.
The theorem follows immediately from Lemma 3.2 and the criterion provided in Bethke, Klop, and de Vrijer [4]. ∎
We point out again that Theorem 3.3 is for the coding of $\mathcal{K}_{2}$ from Shafer and Terwijn [22]. This coding is equivalent to the original coding from Kleene and Vesley [12] in the sense that there are weak embeddings back and forth, as proved in Golov and Terwijn [10]. It does not automatically follow from this that Theorem 3.3 also holds for the original coding. That it does will follow from Theorem 4.4.
We now give a second, much quicker, argument for the strong completability of $\mathcal{K}_{2}$. Recall that $\mathcal{B}$ denotes van Oosten’s model, obtained by extending $\mathcal{K}_{2}$ to partial functions. The embedding $\mathcal{K}_{2}\hookrightarrow\mathcal{B}$ is strong, since it is just the inclusion, and we can choose the combinators $s$ and $k$ of $\mathcal{K}_{2}$ so that they also work for $\mathcal{B}$. Since $\mathcal{B}$ is a total combinatory algebra, it follows immediately that $\mathcal{K}_{2}$ is strongly completable. The same remarks apply for the embedding $\mathcal{K}_{2}^{\mathrm{eff}}\hookrightarrow\mathcal{B}^{\mathrm{eff}}$ of the effective versions of these pcas.
4. $\mathcal{K}_{2}$ generalized
We will make use of the following generalization of $\mathcal{K}_{2}$, taken from van Oosten [19, 1.4.4].
The basic idea underlying $\mathcal{K}_{2}$ is that every partial continuous $f:\omega^{\omega}\to\omega^{\omega}$ can be coded by a real $\alpha\in\omega^{\omega}$. Namely, $f$ can be defined by its modulus of continuity $\hat{f}:\omega^{<\omega}\to\omega^{<\omega}$, which is essentially a function $\alpha:\omega\to\omega$ by virtue of $\omega^{<\omega}=\omega$ (as cardinals). This observation allows use to define a partial application $\alpha\cdot\beta$ on $\omega^{\omega}$ by applying the partial continuous function coded by $\alpha$ to $\beta$.
Now suppose that $\kappa$ is a cardinal such that $\kappa^{<\kappa}=\kappa$. Then the same ideas apply to show that every partial continuous $f:\kappa^{\kappa}\to\kappa^{\kappa}$ can be coded by an element $\alpha\in\kappa^{\kappa}$. So again we can define a partial application $\alpha\cdot\beta$ on $\kappa^{\kappa}$ by interpreting $\alpha$ as a partial continuous function on $\kappa^{\kappa}$. The proof that this yields a pca proceeds along the same lines as the proof for $\mathcal{K}_{2}$. This pca is denoted by $\mathcal{K}_{2}^{\kappa}$.
For the proofs below (especially that of Theorem 5.1), we need to be more specific about how $\alpha\in\kappa^{\kappa}$ codes a partial continuous function. Unfortunately, for large cardinals we have no machine model available, [^2] This could be remedied to some extent by using tools from higher computability theory, but also these have a limit, and the cardinals we need here may be even larger. so that we have to use the less user friendly but more concrete coding of Kleene, which we now specify.
Following Kunen [14, p34], we define $\kappa^{<\kappa}=\bigcup\{\kappa^{\lambda}:\lambda<\kappa\}$. For a function $\sigma\in\kappa^{<\kappa}$, we let $\mathrm{dom}(\sigma)$ denote the largest $\lambda<\kappa$ on which it is defined. By assuming $\kappa^{<\kappa}=\kappa$, we have a bijective coding function $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. For $n\in\kappa$ and $\sigma$ an element of either $\kappa^{<\kappa}$ or $\kappa^{\kappa}$, we let $n\mbox{}^{\smallfrown}\sigma$ denote the function defined by
$(n\mbox{}^{\smallfrown}\sigma)(0)$ $=n$
$(n\mbox{}^{\smallfrown}\sigma)(m+1)$ $=\sigma(m)\text{ if }m<\omega$
$(n\mbox{}^{\smallfrown}\sigma)(m)$ $=\sigma(m)\text{ if }\omega\leqslant m<\kappa.$
For $\alpha\in\kappa^{\kappa}$ and $n\in\kappa$, we let $\alpha\mbox{\raisebox{1.42262pt}{$\upharpoonright$}}n$ denote the restriction of $\alpha$ to elements smaller than $n$.
Define a topology on $\kappa^{\kappa}$ that has as basic open sets
$U_{\sigma}=\big{\{}\alpha\in\kappa^{\kappa}\mid\forall n\in\mathrm{dom}(\sigma)\;\alpha(n)=\sigma(n)\big{\}}$
with $\sigma\in\kappa^{<\kappa}$. Then every $\alpha\in\kappa^{\kappa}$ defines a partial continuous map $F_{\alpha}:\kappa^{\kappa}\to\kappa$ as follows:
[…]
If there are no such $n$ and $m$ we let $F_{\alpha}(\beta)$ be undefined.
Definition 4.1.
$\mathcal{K}_{2}^{\kappa}$ is the pca on $\kappa^{\kappa}$ with application $\alpha\cdot\beta$ defined by
$(\alpha\cdot\beta)(n)=F_{\alpha}(n\mbox{}^{\smallfrown}\beta),$
where it is understood that $\alpha\cdot\beta\!\downarrow$ if $(\alpha\cdot\beta)(n)\!\downarrow$ for every $n\in\kappa$.
Lemma 4.2.
Suppose $F:\kappa^{\kappa}\times\kappa^{\kappa}\to\kappa^{\kappa}$ is continuous. Then there exists $\phi\in\kappa^{\kappa}$ such that $\phi\alpha\beta=F(\alpha,\beta)$ for all $\alpha,\beta\in\kappa^{\kappa}$.
Proof.
This statement can be proved in the same way as Lemma 1.4.1 in van Oosten [19]. The easiest way to do this is by following the proof of Lemma 12.2.2 in Longley and Normann [16]. ∎
Lemma 4.2 is useful in showing that $\mathcal{K}_{2}^{\kappa}$ is a pca. The existence of a combinator $k\in\kappa^{\kappa}$ such that $k\alpha\beta=\alpha$ for all $\alpha,\beta\in\kappa^{\kappa}$ follows immediately from the lemma. We discuss the existence of the combinator $s$ in the proof of Lemma 4.3 below. This lemma generalizes Lemma 3.2 to $\mathcal{K}_{2}^{\kappa}$. Note that there is no machine model available for $\mathcal{K}_{2}^{\kappa}$, so that we are forced to use the coding above. The resulting proof unfortunately has little beauty in it, least of all “cold and austere”, but it does give us what we need, namely Theorem 4.4.
Lemma 4.3.
$\mathcal{K}_{2}^{\kappa}$ has unique hnfs.
Proof.
The existence of a combinator $s\in\kappa^{\kappa}$ such that $s\alpha\beta$ is always defined and $s\alpha\beta\gamma\simeq\alpha\gamma(\beta\gamma)$ for all $\alpha,\beta,\gamma\in\kappa^{\kappa}$ follows from Lemma 4.2 if we can define a (total) continuous binary function $F$ such that
$F(\alpha,\beta)\gamma\simeq\alpha\gamma(\beta\gamma)$ (8)
for all $\alpha,\beta,\gamma$. Writing out (8) using Definition 4.1 shows exactly how to define $F$, and that it is indeed total and continuous in $\alpha$ and $\beta$.
Now if we can also make $F$ injective, it will follow that $s$ such that $s\alpha\beta=F(\alpha,\beta)$, is also injective, and hence satisfies Barendregt’s axiom (5). However, the definition of $F$ obtained from unwinding (8) does not automatically give that $F$ is injective, so we need to do a bit of extra work for this. We show that, starting with any $F$ satisfying (8), we can convert $F$ to a function that is in addition injective. The idea is that if $(F(\alpha,\beta)\gamma)(n)$ is defined for just any $\gamma\in\kappa^{\kappa}$ and $n\in\kappa$, then the coding of Definition 4.1 leaves ample room to code $\alpha$ and $\beta$ in the values of $F(\alpha,\beta)$ that do not matter. The only trouble is when there are no defined values at all, in which case we have to resort to another solution.
Case 1. Suppose $(F(\alpha,\beta)\gamma)(n)$ is defined for some $\gamma\in\kappa^{\kappa}$ and $n\in\kappa$. By Definition 4.1 this means that $F_{F(\alpha,\beta)}(n\mbox{}^{\smallfrown}\gamma)$ is defined. This definition uses only an initial segment $\gamma\mbox{\raisebox{1.42262pt}{$\upharpoonright$}}m$ for some $m\in\kappa$, namely the values $F(\alpha,\beta)(\langle n\mbox{}^{\smallfrown}\gamma\mbox{\raisebox{1.42262pt}{$\upharpoonright$}}m^{\prime}\rangle)$ with $m^{\prime}\leqslant m$. In particular, the higher values $F(\alpha,\beta)(\langle n\mbox{}^{\smallfrown}\gamma\mbox{\raisebox{1.42262pt}{$\upharpoonright$}}m^{\prime}\rangle)$ for $m^{\prime}>m$ are irrelevant, and we are free to alter them as we like, without affecting the effect of $F$. Now we use these values to code the sequences $\alpha$ and $\beta$. Using a bijection $\kappa+\kappa\leftrightarrow\kappa$, we can code them as one sequence $\alpha\oplus\beta$ of length $\kappa$, by redefining the values $F(\alpha,\beta)(\langle n\mbox{}^{\smallfrown}\gamma\mbox{\raisebox{1.42262pt}{$\upharpoonright$}}m^{\prime}\rangle)$ with $m^{\prime}>m$.
Case 2. $(F(\alpha,\beta)\gamma)(n)$ is undefined for all $\gamma\in\kappa^{\kappa}$ and $n\in\kappa$. In this case we have by Definition 4.1 that $F(\alpha,\beta)(n)$ is equal to 00 or a limit for all $n\in\kappa$. The effect of $F(\alpha,\beta)$ does not change if we change the value from one limit or 00 to another. W.l.o.g. $\kappa>\omega$, since for $\kappa=\omega$ we already have a proof (Lemma 3.2). If $\kappa>\omega$ we have the two values $0,\omega<\kappa$ that we can use for coding $\alpha$ and $\beta$. Again we can code these as one sequence $\alpha\oplus\beta\in\kappa^{\kappa}$. Using a bijection $\kappa^{\kappa}\leftrightarrow 2^{\kappa}$ we may assume $\alpha\oplus\beta\in 2^{\kappa}$, and code it using the two values 00 and $\omega$ by
$F(\alpha,\beta)(n)=\begin{cases}0&\text{if $(\alpha\oplus\beta)(n)=0$},\\ \omega&\text{if $(\alpha\oplus\beta)(n)=1$}.\end{cases}$
Now since the adapted version of $F$ still has to be continuous, we cannot make the case distinction above separately for every $\alpha$ and $\beta$: For some $\alpha$ and $\beta$ where case 1 applies, there may be close $\alpha^{\prime}$ and $\beta^{\prime}$ where case 2 applies, and since the coding for $\alpha^{\prime}$ and $\beta^{\prime}$ starts from the beginning, we have no choice but to do the same for $\alpha$ and $\beta$. So we define the adaptation of $F(\alpha,\beta)(\langle\sigma\rangle))$, with $\sigma\in\kappa^{<\kappa}$, by switching between the two cases dynamically. As long as we have not found a point yet where case 1 applies, we follow the coding of case 2, and otherwise we switch to the coding of case 1. Specifically, if for every $\sigma^{\prime}\sqsubseteq\sigma$, $F(\alpha,\beta)(\langle\sigma^{\prime}\rangle))$ is 00 or a limit, we apply the coding of case 2. Otherwise, if $\sigma^{\prime}\sqsubseteq\sigma$ is the minimal initial such that $F(\alpha,\beta)(\langle\sigma^{\prime}\rangle))$ is a successor, we apply from that point onwards the coding of case 1.
This concludes the proof that we can make $F$ injective without changing its behavior. As we already noted, the resulting combinator $s$ then satisfies Barendregt’s axiom (5). In particular, all the hnfs $s\alpha\beta$ are unique. Since $\alpha$ can be retrieved from $s\alpha\beta$ for any $\beta$, it follows that $s\alpha$ is also injective, and hence the hnfs $s\alpha$ are also unique. It follows from the properties of $k$ that these are also different from all $k\alpha$, so that indeed all hnfs are different. ∎
Theorem 4.4.
Every $\mathcal{K}_{2}^{\kappa}$ is strongly completable.
Proof.
This is again immediate from Lemma 4.3 and the criterion provided in Bethke, Klop, and de Vrijer [4]. ∎
For $\mathcal{K}_{2}$ (i.e. for $\kappa=\omega$), we gave two arguments for its strong completability in Section 3. Apart from the argument using Lemma 3.2 about unique hnfs we had an easier option using van Oosten’s model $\mathcal{B}$. It would be interesting to generalize this to $\mathcal{B}^{\kappa}$ for all $\kappa$ with $\kappa^{<\kappa}=\kappa$ as above. It is not immediately clear how to define $\mathcal{B}^{\kappa}$. Simply restricting $\mathcal{K}_{2}^{\kappa}$ to partial functions does not work, as this makes use of the machine model for $\mathcal{K}_{2}$ that is not available for $\mathcal{K}_{2}^{\kappa}$. Also, the coding from Definition 4.1 is not suitable for $\mathcal{B}$, as is discussed in Longley and Normann [16, p77], because this coding does not allow to jump over undefined values.
The definition of $\mathcal{K}_{2}^{\kappa}$ uses the property $\kappa^{<\kappa}=\kappa$, needed for the coding $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. It is known that inaccessible cardinals have this property, cf. Jech [11, p50], but the existence of inaccessibles cannot be proved in $\mathrm{ZFC}$ (cf. Kunen [14]). For $\kappa=2^{\omega}$, the property $\kappa^{<\kappa}=\kappa$ is independent. [^3] One can easily check that the statement is true under CH and false e.g. when $2^{\omega}=\omega_{2}$ and $2^{\omega_{1}}=\omega_{3}$. To guarantee existence of $\mathcal{K}_{2}^{\kappa}$ for arbitrary large $\kappa$ we can for example use the generalized continuum hypothesis $\mathrm{GCH}$, which says that $2^{\omega_{\alpha}}=\omega_{\alpha+1}$ for every ordinal $\alpha$. (This is consistent by Gödels famous result that $\mathrm{GCH}$ holds in $\mathrm{L}$.) Assuming $\mathrm{GCH}$ we have
$\omega_{\alpha+1}^{<\omega_{\alpha+1}}=(2^{\omega_{\alpha}})^{\omega_{\alpha}}=2^{\omega_{\alpha}}=\omega_{\alpha+1}$ (9)
so that $\kappa=\omega_{\alpha+1}$ satisfies the hypothesis we need for $\mathcal{K}_{2}^{\kappa}$.
5. Embeddings into $\mathcal{K}_{2}^{\kappa}$
Having seen that every countable pca weakly embeds into $\mathcal{K}_{2}$, we want to extend this result to larger cardinalities. Now we do have the large pcas $\mathcal{K}_{2}^{\kappa}$, which is a pca on $\kappa^{\kappa}$, but as discussed in Section 4 the definition of this is contingent on the condition $\kappa^{<\kappa}=\kappa$, which implies that for $\kappa>\omega$ at best we know that their existence is consistent with $\mathrm{ZFC}$ (using $\mathrm{GCH}$).
Theorem 5.1.
Assuming $\mathrm{GCH}$, for every pca $\mathcal{A}$ there exists a cardinal $\kappa$ such that $\mathcal{A}$ weakly embeds into $\mathcal{K}_{2}^{\kappa}$.
Proof.
Assume $\mathrm{GCH}$, and let $\mathcal{A}$ be any pca. Let $\alpha$ be an ordinal such that $|\mathcal{A}|\leqslant\omega_{\alpha+1}$, and let $\kappa=\omega_{\alpha+1}$. By (9) we have $\kappa^{<\kappa}=\kappa$, and hence the pca $\mathcal{K}_{2}^{\kappa}$ exists. We prove that $\mathcal{A}$ weakly embeds into $\mathcal{K}_{2}^{\kappa}$.
The proof from [10] that every countable pca $\mathcal{A}$ embeds into $\mathcal{K}_{2}$ uses the recursion theorem, which we do not have available for large cardinalities, so in the proof below we have to avoid this. But we can still use the idea that because $|\mathcal{A}|\leqslant\kappa$, the graph of the application operator in $\mathcal{A}$ can be coded by an element of $\kappa^{\kappa}$.
Since $|\mathcal{A}|\leqslant\kappa$, we can fix a numbering $\gamma:\kappa\to\mathcal{A}$. If $\gamma(n)=a$ we call $n$ a code of $a$. We call $n\in\kappa$ a minimal code if for all $m<n$, $\gamma(m)\neq a$.
To obtain an embedding $f:\mathcal{A}\hookrightarrow\mathcal{K}_{2}^{\kappa}$ we define $f(a)\in\kappa^{\kappa}$ for every $a\in\mathcal{A}$ such that
$\mathcal{A}\models a\cdot b\!\downarrow=c\;\Longrightarrow\;\mathcal{K}_{2}^{\kappa}\models f(a)\cdot f(b)\!\downarrow=f(c).$ (10)
Define a subset $L\subseteq\kappa$ by
$L_{0}$ =$=\{0\}$
$L_{i+1}$ =$=\{\langle m\mbox{}^{\smallfrown}z\rangle\mid m\in L_{i}\wedge z\in\kappa\}$
$L$ $=\bigcup_{i\in\omega}L_{i}.$
[…]
Referring to the coding before Definition 4.1, we define $f(a)$ by:
  1. (i)
    Case $L_{0}$. $f(a)(0)=w_{0}=$ minimal code of $a$.
  2. (ii)
    Case $L_{1}$. $f(a)(\langle 0\mbox{}^{\smallfrown}z_{0}\rangle)=$ minimal code +1 of $\gamma(w_{0})\cdot\gamma(z_{0})\in\mathcal{A}$ if this is defined, and 00 otherwise.
  3. (iii)
    Case $L_{i}$. $f(a)(\langle n\mbox{}^{\smallfrown}z_{0}\rangle)$ for $n\in L_{i-1}$, $i\geqslant 1$, is defined as follows. Note that $n$ is of the form
    $n=\langle\langle\langle\ldots\langle\langle 0\mbox{}^{\smallfrown}z_{i-1}\rangle\mbox{}^{\smallfrown}z_{i-2}\rangle\mbox{}^{\smallfrown}\ldots\rangle\mbox{}^{\smallfrown}z_{1}\rangle\mbox{}^{\smallfrown}z_{0}\rangle.$ (11)
    Define
[…]
    $\vdots$
[…]
    Define $f(a)(\langle n\mbox{}^{\smallfrown}z_{0}\rangle)=w_{i}+i$ provided that all applications in the definition of $w_{1},\ldots,w_{i}$ are defined, and $f(a)(\langle n\mbox{}^{\smallfrown}z_{0}\rangle)=0$ otherwise.
  4. (iv)
    $f(a)(n)=0$ for all $n\notin L$.
Note that with this definition, the graph $\{(a,b,c)\mid\mathcal{A}\models a\cdot b\!\downarrow=c\}$ is coded into $f(a)$ for every $a$. Now we should check that (10) holds. Suppose that $a\cdot b\!\downarrow=c$ in $\mathcal{A}$. We check that Definition 4.1 gives $(f(a)\cdot f(b))(n)=f(c)(n)$ for every $n\in\kappa$. Suppose that $u,v,w$ are minimal codes of $a,b,c$, respectively.
Case $n=0$. By (6) we have $(f(a)\cdot f(b))(0)=F_{f(a)}(0\mbox{}^{\smallfrown}f(b))=f(a)(\langle 0\mbox{}^{\smallfrown}v\rangle)-1$. We use here that $f(a)(\langle\emptyset\rangle)=0$ (because we have assumed $\langle\emptyset\rangle\notin L$) and $f(a)(\langle 0\rangle)=0$ (because $\langle 0\rangle\notin L$), so that (7) is fulfilled. Now by (ii) we have $f(a)(\langle 0\mbox{}^{\smallfrown}v\rangle)=$ minimal code $+1$ of $\gamma(u)\cdot\gamma(v)=c$, which is $w+1$. By (i) we have $f(c)(0)=w$, so that indeed $(f(a)\cdot f(b))(0)=f(c)(0)$.
Case $n\in L_{i}$, $i\geqslant 1$. Suppose $n$ is of the form (11). By (iii) we have $f(c)(n)=w_{i}+i$, starting with $w_{1}$ a minimal code of $\gamma(w)\cdot\gamma(z_{0})$, where $w$ is now the minimal code of $c$.
This should equal $(f(a)\cdot f(b))(n)=F_{f(a)}(n\mbox{}^{\smallfrown}f(b))=f(a)(\langle n\mbox{}^{\smallfrown}v\rangle)-1$. Indeed, $f(a)(\langle n\mbox{}^{\smallfrown}v\rangle)=w_{i}+i+1$, where we start with $w$ a minimal code of $\gamma(u)\cdot\gamma(v)$, $w_{1}$ a minimal code of $\gamma(w)\cdot\gamma(z_{0})$, etc. So we see that $(f(a)\cdot f(b))(n)=f(c)(n)$, provided that also (7) is fulfilled.
For $f(a)(\langle n\mbox{}^{\smallfrown}v\rangle)$ to produce the desired value, we need for (7) that both $f(a)(\langle n\rangle)$ and $f(a)(\langle\emptyset\rangle)$ are 00. Again, by assumption, $\langle\emptyset\rangle\notin L$, so $f(a)(\langle\emptyset\rangle)=0$. For every $n\in L$ we have $\langle n\rangle\notin L$, so indeed also $f(a)(\langle n\rangle)=0$ by (iv). ∎
6. Concluding remarks on completions
Asperti and Ciabattoni [1] proved that if Barendregt’s axiom (5) holds in a pca then it has a weak completion. (See also the discussion about this in Bethke et al. [5, p501].) By Engeler [6] (see footnote 1 above), the condition about Barendregt’s axiom is not needed. Note that Barendregt’s axiom does not hold for the example from Bethke et al. [5] of a pca without strong completions. Indeed, on p486 loc. cit. we have $a$ and $b$ distinct such that $s(sk)a=s(sk)b$, which violates Barendregt’s axiom. But since the example constructed in [5] is countable, and every countable pca weakly embeds into $\mathcal{K}_{2}$, it also follows from the strong completability of $\mathcal{K}_{2}$ that it has a weak completion.
Theorem 5.1 shows that, under GCH, every pca weakly embeds into a $\mathcal{K}_{2}^{\kappa}$ for suitably large $\kappa$. Since each $\mathcal{K}_{2}^{\kappa}$ is strongly completable, this has as a corollary that every pca has a weak completion (still assuming GCH). The latter statement follows unconditionally from Engeler [6]. GCH is needed in the proof of Theorem 5.1 to guarantee the existence of large cardinals $\kappa$ with $\kappa^{<\kappa}=\kappa$. However, we can get the corollary about completions unconditionally by using a binary version of $\mathcal{K}_{2}^{\kappa}$ defined on $2^{\kappa}$ rather than $\kappa^{\kappa}$. Namely, for the binary version we need the condition $2^{<\kappa}=\kappa$, and ZFC unconditionally proves the existence of arbitrary large $\kappa$ with this property. [^5] Namely, $2^{<\kappa}=\kappa$ holds for any $\kappa$ that is a strong limit, i.e. such that $\lambda<\kappa$ implies $2^{\lambda}<\kappa$, and the strong limit cardinals form a proper class, cf. Jech [11, p50]. This gives an alternative proof of Engeler’s result that every pca is weakly completable.
Note that the condition $\kappa^{<\kappa}=\kappa$ is needed in the definition of $\mathcal{K}_{2}^{\kappa}$ as the analog of the coding of finite sequences $\omega^{<\omega}=\omega$ that is used in the definition of $\mathcal{K}_{2}$. Coding of finite sets also plays a role in the definition of Scott’s graph model $\mathcal{G}$ mentioned above. It is interesting to see how Engeler avoids this condition by explicitly putting the finite sets into his definition of a graph algebra $G(A)$. This has as advantage that no coding is needed. On the negative side, the elements of $G(A)$ are somewhat contrived, and much more complicated than the elements of $\omega^{\omega}$ or $\kappa^{\kappa}$.
Acknowledgments
We thank Henk Barendregt for pointing out that it follows from Engeler [6] that every pca has a weak completion, and for sharing the supplementary note Engeler [7]. We also acknowledge the helpful remarks of a referee that corrected an error in the proof of Lemma 4.3.